OpenStudy (anonymous):

5(2x – 3) = 5 Part A: How many solutions does this equation have? Part B: What are the solutions to this equation? Show your work.

pooja195 (pooja195):

Start by distributing 5(2x – 3)

OpenStudy (anonymous):

5*2x-15? @pooja195

pooja195 (pooja195):

5(2x−3)=5 (5)(2x)+(5)(−3)=5(Distribute)

OpenStudy (anonymous):

i dont understand @pooja195

pooja195 (pooja195):

like that

OpenStudy (anonymous):

yeah i know that so it would be 10x-15? @pooja195

pooja195 (pooja195):

yep :) 10x-15=5 15+5=?

OpenStudy (anonymous):

20? @pooja195

pooja195 (pooja195):

good 20/10?

OpenStudy (anonymous):

2 @pooja195

pooja195 (pooja195):

good so the answer is?

OpenStudy (anonymous):

x=2? @pooja195

pooja195 (pooja195):

yes

OpenStudy (anonymous):

okay, one more? @pooja195

pooja195 (pooja195):

ok :)

OpenStudy (anonymous):

An equation is shown below: 9(3x – 16) + 15 = 6x – 24 Part A: Write the steps you will use to solve the equation, and explain each step. (6 points) Part B: What value of x makes the equation true? (4 points) @pooja195

pooja195 (pooja195):

distribute 9(3x – 16)

OpenStudy (anonymous):

27x-144? @pooja195

pooja195 (pooja195):

good 27x-144+15=6x – 24 combine like terms

OpenStudy (anonymous):

21x-105? @pooja195

pooja195 (pooja195):

no -144+15=?

OpenStudy (anonymous):

21x-129*? @pooja195

pooja195 (pooja195):

Step 1: Simplify both sides of the equation. 27x−144+15=6x−24 Simplify: (Hide steps) Let's simplify step-by-step. 27x−144+15=6x−24 27x+−144+15=6x+−24 (27x)+(−144+15)=6x−24(Combine Like Terms) 27x+−129=6x−24 27x−129=6x−24 Step 2: Subtract 6x from both sides. 27x−129−6x=6x−24−6x 21x−129=−24 Step 3: Add 129 to both sides. 21x−129+129=−24+129 21x=105 Step 4: Divide both sides by 21.

pooja195 (pooja195):

Credits: http://www.mathpapa.com/algebra-calculator.html